Change the equation to quadratic form and solve it by the most efficient method.

45/b - 16 = - 16/b+1

Answer by stanbon(75887)

45/b - 16 = - 16/b+1
---
(45/b)+(16/b) = 17
---
(45+16)/b = 17
----
61/b = 17
b = 61/17
----
Cheers,
Stan H.